A factory manufacturing light-emitting diode (LED) bulb claims that their light bulbs last for 50,000 hours on the average. To confirm if this claim was valid, a quality control manager got a sample of 50 LED bulbs and obtained a mean lifespan of 40,000 hours. The standard deviation of the manufacturing process is 20,000 Do you think the claim of the manufacturer is valid at the 5% level of significance? 


1
Expert's answer
2021-06-29T12:43:35-0400

"\bar{x}=40000 \\

s=20000 \\

n=50 \\

df=n-1=49 \\

α=0.05 \\

H_0: \mu=50000 \\

H_1: \mu ≠50000 \\

t_{crit}=2.095"

The rejection region for this two-tailed test is R={t:|t|>2.0095}

Test-statistic:

"t = \frac{\bar{x}- \mu}{s/ \sqrt{n}} \\

t = \frac{40000-50000}{20000 / \sqrt{50}} = -3.535 \\

|t|=3.535>2.0095"

The null hypothesis is rejected.

There is enough evidence to claim that the population mean μ is different than 50000, at the 0.05 significance level.